OpenStudy (anonymous):

Which of the following constants can be added to x2 - 3x to form a perfect square trinomial? 3/4 1 1/2 2 1/4

OpenStudy (anonymous):

"Which of the following constants can be added to x² – 3x to form a perfect square trinomial? 3/4, or 1(1/2), or 2(1/4) A: As the offered answers have been presented? None of them! Now if by: • "1(1/2)" you meant "1½" • "2 (1/4)" you meant "2¼" Ahhhh! That changes things A LOT! IF ––– that is what you meant THEN ––– 2¼ is correct ... and here is why: x² – 3x + (?)² = [ x – ?]² = x² – 2(x)(?) + ?² "?" must be such that (–2)(x)(?) = –3x ➞ (–2)(?) = –3 ➞ ? = –3/(–2) = 3/2 So (3/2)² = 9/4 = 2¼ Hope this helps!

OpenStudy (anonymous):

9/4